### v2

@@ 2 Related Work
Many works are available related to architectures for healthcare IoT targeting various aspects. This includes, analysis of IoT architectures for healthcare applications [22], design patterns for healthcare IoT [18], architecture for intelligent IoT-based healthcare systems [10], architectural design decisions for developing digital twins of IoT systems [16], a tool for modeling IoT architectures [30], health monitoring architecture for IoT systems [6], an architecture for IoT-based remote patient monitoring [2], a requirements-based healthcare IoT architecture [15], distributed IoT architecture [3], health data sharing architecture [23], analyze architecture for healthcare IoT system [25], architecture for an IoT-based healthcare monitoring system [19], and architecture for blockchain-driven healthcare IoT systems [31]. In contrast to the aforementioned works, our work focuses on developing a test infrastructure architecture for healthcare IoT applications.
Several works are also available targeting architecture-based testing. This includes, analysis of architecture’s function in software testing [7], architecture-based test criteria [14], regression testing with software architectures [20], architecture-driven integration testing [21], an architecture for analyzing fault tolerance [8], and reliability assessment using architecture models [9]. Compared to the above-mentioned works, our work offers a distinct contribution by presenting a test infrastructure architecture for healthcare IoT applications, which is designed to facilitate automated system-level testing of such applications.
@@ 3.1 HITA Components
[…]
@@
[…]
@@ 3.2 Quality Attributes
3.2.9 Portability
is an additional feature of HITA. The architecture followed by TSGen and DTGen components can work on local machines when testing offline and remotely in different cloud environments.
@@ 4 Evaluation
    In this RQ, we analyze the scalability of operating 100 DTs in different batch sizes (i.e., 10, 20, .., 100) during test execution.
@@ 4.1 DTGen Component Implementation
To generate model-based DT, we follow the approach presented in [26], which is briefly described below. In the first step, we model the structural and behavioral aspects of a medical device for which we need to create DTs. This requires creating a domain model of a medical device to capture device concepts and properties, specifying constraints on device properties using Object Constraint Language (OCL), and modeling the device behavior in the form of state machines. Creating models and constraints in the first step is manual. The next steps are automated. In the second step, we create an instance model of the domain model using the input device property values (in JSON format) and validate the instance model using OCL constraints. In the third step, we create a device state machine as the owned behavior of the device instance model. Finally, we make a device executable model to operate DT simulating the device.
@@ 4.4 Results and Discussion
RQ1 Result
@@
RQ2 Result
@@
Table 2 and 3 present the time cost involved in various steps of creating MBDT and MLDT, respectively. To generate MBDT, modeling domain, constraints, and behavior requires significant time, approximately 1-2 hours, which can vary depending on the modelers’ experience. Providing input configurations may take 2-5 minutes. It is important to note that generating one or more MBDTs is an automated step and only takes time in milliseconds. The time required to generate one MBDT and 100 MBDTs is, on average, $\approx$108 and $\approx$229 milliseconds, respectively.
@@
RQ3 Result
@@ 6 Conclusion and Future Work
In this paper, we presented real-world architectural work in collaboration with Oslo City’s healthcare department. We introduced HITA – a test infrastructure architecture to facilitate automated system-level testing of healthcare IoT applications, with design considerations aligned to Oslo City’s healthcare department requirements. We evaluated HITA’s DTGen component by creating DTs of a Medido medicine dispenser using model-based and ML approaches. Our evaluation focused on analyzing the fidelity of DT created using both approaches, assessing the scalability of the DTGen component when operating 100 DTs in different batch sizes, and the time cost involved in creating DTs. Results show that the fidelity of DTs created using model-based and ML approaches is 94% and 95%, respectively. Results with 100 DTs also show that the DTGen component is scalable. Moreover, results indicate that DTs generated with the ML approach have fewer manual steps and higher time costs compared to model-based DTs. Finally, we presented experience and lessons learned based on experiments conducted with work products of HITA that are valuable for industry practitioners working in a similar domain. Our lessons learned are generalizable to various IoT-based systems such as activity/fitness trackers, smart homes, and smart security systems.

### v3

@@ 2 Related Work
Many works are available related to architectures for healthcare IoT targeting various aspects. This includes; analysis of IoT architectures for healthcare applications [21], design patterns for healthcare IoT [18], architecture for intelligent IoT-based healthcare systems [10], architectural design decisions for developing digital twins of IoT systems [16], a tool for modeling IoT architectures [30], health monitoring architecture for IoT systems [6], an architecture for IoT-based remote patient monitoring [2], a requirements-based healthcare IoT architecture [15], distributed IoT architecture [3], health data sharing architecture [22], analyze architecture for healthcare IoT system [24], architecture for an IoT-based healthcare monitoring system [19], and architecture for blockchain-driven healthcare IoT systems [31]. In contrast to the aforementioned works, our work focuses on developing a test infrastructure architecture for healthcare IoT applications.
Several works are also available targeting architecture-based testing. This includes; analysis of architecture’s function in software testing [7], architecture-based test criteria [14], architecture-driven integration testing [20], an architecture for analyzing fault tolerance [8], and reliability assessment using architecture models [9]. Compared to the above-mentioned works, our work offers a distinct contribution by presenting a test infrastructure architecture for healthcare IoT applications, which is designed to facilitate automated system-level testing of such applications.
@@ 3.1 HITA Components
For integration with medical devices, HITA utilizes the concept of digital twins to create a virtual representation of physical devices. Each medical device from a different vendor is connected to a server with several APIs for integration (as shown in the right-bottom of Figure 1). The architecture for creating physical medical devices DTs consists of one DT Server with APIs (e.g., APIs DT-D1) specific to a certain type of DTs representing a particular device (e.g., DTs-D1). APIs need to be developed following REST architecture [12] to allow easy integration with SUT. DTs of medical devices can be generated with commonly used approaches [32] like model-based approach or ML approach. In the case of multiple versions of a medical device, a separate DT for each device variant is required to be generated. For the communication between the DT Server and DTs, JSON data interchange format is used. In case medical devices support different JSON schema, the Schema Registry can be used to ensure compatibility. The DTGen component also consists of Data Persistence to preserve the state of DTs among various requests. The APIs of DTs are used to integrate DTs with SUT and physical medical devices. During testing, the DTs act as middle-ware between SUT and physical medical devices. DTs handle all communication traffic from SUT and communicate (via HTTP) with their physical twins when necessary.
@@
To handle the challenges of integrating third-party applications for testing purposes, HITA’s TSGen component plays an important role. Each third-party application has dedicated servers with APIs for integration. The architecture for test stubs creation consists of one TS Server with APIs (e.g., APIs TS1) simulating the behavior of various applications (e.g., App 1). The APIs for each test stub must be developed according to REST architecture [12] for easy integration with SUT. Test stubs play a key role in replicating the functionality of third-party applications. For APIs requiring data (e.g., health data), the architecture includes an artificial data store with multiple databases corresponding to APIs representing different applications. The data manipulation is performed using query language compliant with the database type.
@@ 3.2 Quality Attributes
3.2.9 Portability.
It is an additional feature of HITA. The architecture followed by TSGen and DTGen components can work on local machines when testing offline and remotely in different cloud environments.
@@ 4 Evaluation
    In this RQ, we assess the scalability of HITA by examining its ability to manage the operation of 100 DTs across various batch sizes (i.e., 10, 20, .., 100) during test execution.
@@ 4.1 DTGen Component Implementation
To generate model-based DT, we follow the approach presented in [26], which is briefly described below. In the first step, we model the structural and behavioral aspects of a medical device for which we need to create DTs. This requires creating a domain model of a medical device to capture device concepts and properties, specifying constraints on device properties using Object Constraint Language (OCL), and modeling the device behavior in the form of state machines. Creating models and constraints in the first step is manual. The next steps are automated. In the second step, we create an instance model of the domain model using the input device property values (in JSON format) and validate the instance model using OCL constraints [27]. In the third step, we create a device state machine as the owned behavior of the device instance model. Finally, we make a device executable model to operate DT simulating the device.
@@ 4.4 Results and Discussion
RQ1 Result DTs generated using model-based and ML approaches have fidelity compared to Medido.
@@
RQ2 Result The fidelity of 100 DTs (MBDTs and MLDTs) is consistent in all batch sizes, highlighting that the DTGen component is scalable.
@@
Table 2 and 3 present the estimated time cost involved in various steps of creating MBDT and MLDT, respectively. To generate MBDT, modeling domain, constraints, and behavior requires significant time, approximately 1-2 hours, which can vary depending on the modelers’ experience. Providing input configurations may take 2-5 minutes. It is important to note that generating one or more MBDTs is an automated step and only takes time in milliseconds. The time required to generate one MBDT and 100 MBDTs is, on average, $\approx$108 and $\approx$229 milliseconds, respectively.
@@
RQ3 Result DTs generated using ML require fewer manual steps and more time compared to model-based DT. ML approach with higher time cost generates DTs with higher fidelity compared to model-based DTs.
@@ 6 Conclusion and Future Work
In this paper, we presented real-world architectural work in collaboration with Oslo City’s healthcare department. We introduced HITA – a test infrastructure architecture to facilitate automated system-level testing of healthcare IoT applications, with design considerations aligned to Oslo City’s healthcare department requirements. We evaluated HITA’s DTGen component by creating DTs of a Medido medicine dispenser using model-based and ML approaches. Our evaluation focused on analyzing the fidelity of DT created using both approaches, assessing the scalability of the DTGen component when operating 100 DTs in different batch sizes, and the time cost involved in creating DTs. Results show that the fidelity of DTs created using model-based and ML approaches is 94% and 95%, respectively. Results with 100 DTs also show that the DTGen component is scalable. Moreover, results indicate that DTs generated with the ML approach have fewer manual steps and higher time costs compared to model-based DTs. Finally, we presented experience and lessons learned based on experiments conducted with work products of HITA that are valuable for industry practitioners working in a similar domain. Our architecture, findings, and lessons learned are generalizable to various IoT-based systems such as activity/fitness trackers, smart homes, and smart security systems.